# Defending Code Language Models against Backdoor Attacks with Deceptive Cross-Entropy Loss

### v2

@@ 6.1. Generalization
Generalization to larger models. Our study has assessed the efficacy of DeCE across a spectrum of widely-utilized Code Language Models (CLMs) fewer than 1 billion parameters. Given the remarkable capabilities and complexities of larger models we extend our investigation to encompass three additional CLMs (CodeGeeX (Zheng et al., 2023), CodeLlama (Roziere et al., 2023), and DeepSeekCoder (Guo et al., 2024)) all with more than 1B parameter count.
@@
Table 9. Results in LLMs on Lyra.
@@ 7.2. Backdoor Attack and Defense
Backdoor attacks pose a significant threat to neural network models, targeting the training phase rather than the inference phase, which can be classified into token-based, syntax-based, and semantic-based attacks in NLP. Token-based attacks utilize trigger keywords to generate logical trigger sentences, while syntax-based attacks leverage syntactic triggers. For example, Chen et al. (2021b) enhanced the effectiveness of token-based attacks by introducing semantic preservation trigger generation methods with multiple perturbation levels. Qi et al. (2021b) proposed a method that utilizes these triggers, and they also explored the use of text-style transfer techniques to generate more dynamic backdoor samples. Semantic-based attacks focus on creating backdoor training samples that appear more natural to humans. Chan et al. (2020) utilized an autoencoder to generate these samples, enhancing their authenticity. Among these, token-based attacks demonstrate high attack efficiency but are more susceptible to detection. To overcome this limitation, Chen et al. (2021a) proposed BadPre, a method that bypasses detection by randomly inserting triggers multiple times into the input sequence during deployment. In the realm of programming languages, backdoor implantation has gained attention. Researchers have proposed various strategies, including fixed triggers (Wan et al., 2022), rule-based poisoning (Li et al., 2023b), and language model-guided poisoning (Li et al., 2023c).
In terms of defending against backdoor attacks, most of the studies have focused on models used in NLP. Inference-time defense methods (such as ONION (Qi et al., 2021a)) detect and remove discrete words using language model outputs, while training-time defense methods (such as BKI (Chen and Dai, 2021)) identify and remove potentially poisoned samples during training. Other defense methods (such as Moderate-fitting (Zhu et al., 2022) and In-trust loss (Huang et al., 2021)) involve reducing model capacity and training duration or utilizing specific loss functions. In the context of programming languages, defense strategies involve parsing and identifying potentially uncompilable code as poison samples (Li et al., 2023b).
@@
In our study, we focus on developing defense methods against backdoor attacks. Our defense method leverages the ”early learning” phenomenon observed during the training of CLMs. Our proposed method not only showcases enhanced effectiveness but also exhibits a wider applicability scope when compared with previous defense methodologies.
@@ 8. Conclusion
While DeCE has shown promising results in defending against backdoor attacks, we want to optimize its hyper-parameters in the future, which can improve the defense quality and robustness of DeCE against sophisticated attack strategies. Additionally, we would like to investigate its applicability to more areas of code intelligence, such as code defect detection, code summarization, and other tasks.
@@ Acknowledgements.
The authors are grateful for the valuable feedback from domain experts. This work was partially supported by the National Natural Science Foundation of China (NSFC, No. 61972197 and No. 62372232), the Natural Science Foundation of Jiangsu Province (No. BK20201292), the Collaborative Innovation Center of Novel Software Technology and Industrialization, and the Postgraduate Research. T. Chen is partially supported by an oversea grant from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2022A03) and Birkbeck BEI School Project (EFFECT).

### v3

@@ 6.1. Generalization to classification tasks
6.2. Generalization to larger models.
Our study has assessed the efficacy of DeCE across a spectrum of widely-utilized Code Language Models (CLMs) fewer than 1 billion parameters. Given the remarkable capabilities and complexities of larger models we extend our investigation to encompass three additional CLMs (CodeGeeX (Zheng et al., 2023), CodeLlama (Roziere et al., 2023), and DeepSeekCoder (Guo et al., 2024)) all with more than 1B parameter count.
@@
Table 10. Results on the Lyra dataset across three large models.
@@
6.3. Generalization to larger datasets.
Table 11. Results on the larger code generation datasets.
Model Method 1% (HumanEval) 1% (MBPP) 1% (CodeHarmony-test)
Pass@1 ASR Pass@1 ASR Pass@1 ASR
CodeT5p -220M clean 13.41 - 14.60 - 33.99 -
poisoned 11.59 75.61 13.40 49.00 33.99 32.03
DeCE 14.02 1.22 15.20 0.20 33.33 0.00
CodeGen -350M clean 18.90 - 22.00 - 43.14 -
poisoned 18.29 98.78 21.40 98.20 41.18 98.04
DeCE 19.51 2.44 21.40 2.80 41.83 0.65
CodeLlama -7B clean 32.93 - 31.00 - 50.98 -
poisoned 26.83 63.41 25.20 49.60 50.33 52.94
DeCE 31.71 1.22 30.40 0.40 50.33 0.00
In addition to generalization across different tasks and larger models, it is essential to validate the efficacy of DeCE on larger datasets. To this end, we select CodeHarmony [^2] https://huggingface.co/datasets/Flab-Pruner/CodeHarmony, a large-scale code generation dataset which includes 15,800 training samples, 200 validation samples and 153 test samples. Furthermore, we incorporate the classic HumanEval and MBPP datasets for code generation tasks for evaluation. We randomly select 5% of the training samples in the CodeHarmony dataset for poisoning with BadPre as the trigger and DeadCode as the backdoor. We choose three typical CLMs, i.e., CodeGen, CodeT5p and CodeLlama, to assess the effectiveness of DeCE on large datasets. Moreover, since these datasets come with test cases, we use the Pass@1 metric as the evaluation criterion along with the ASR metric on the poisoned dataset.
The results in Table 11 demonstrate the effectiveness of DeCE in mitigating backdoor attacks on large datasets. We present the results of three different models on three different datasets, including the Clean model trained on the clean training set, the Poisoned model trained on the 10% poisoned training set and the Defense model trained with DeCE on the 10% poisoned training set. All models are trained in 2 epochs with a learning rate of 4e-5. We observe that, when we poison the test set, the ASR of the model trained with DeCE is only 0%-2%, which indicates that DeCE is effective in defending against backdoor attacks on all three datasets. Notably, DeCE maintains a high Pass@1 score on the clean test set, indicating that it does not compromise the model’s performance on clean data.
@@ 7.2. Backdoor Attack
Backdoor attacks pose a significant threat to neural network models, targeting the training phase rather than the inference phase, which can be classified into token-based, syntax-based, and semantic-based attacks in NLP. Token-based attacks utilize trigger keywords to generate logical trigger sentences, while syntax-based attacks leverage syntactic triggers. For example, Chen et al. (2021b) enhanced the effectiveness of token-based attacks by introducing semantic preservation trigger generation methods with multiple perturbation levels. Qi et al. (2021b) proposed a method that utilizes these triggers, and they also explored the use of text-style transfer techniques to generate more dynamic backdoor samples. Semantic-based attacks focus on creating backdoor training samples that appear more natural to humans. Chan et al. (2020) utilized an autoencoder to generate these samples, enhancing their authenticity. Among these, token-based attacks demonstrate high attack efficiency but are more susceptible to detection. To overcome this limitation, Chen et al. (2021a) proposed BadPre, a method that bypasses detection by randomly inserting triggers multiple times into the input sequence during deployment. In the realm of programming languages, backdoor implantation has gained attention. Researchers have proposed various strategies, including fixed triggers (Wan et al., 2022), rule-based poisoning (Li et al., 2023b), and language model-guided poisoning (Li et al., 2023c). For instance, Cotroneo et al. (2024) proposed a data poisoning attack to assess the security of code generators by injecting software vulnerabilities to the training data. Sun et al. (2023a) proposed a stealthy backdoor attack BADCODE against neural code search models by modifying variable/function names.
7.3. Backdoor Defense
Most studies defending against backdoor attacks have focused on models used in NLP.
Active Defense. Active defense methods aim to detect and remove backdoor samples defore or during the training phase. For example, Chen and Dai (2021) proposed Backdoor Keyword Identification (BKI), which identifies and removes potential poisoned samples during the training process. Zhu et al. (2022) proposed Moderate-fitting, which defends against backdoor attacks by reducing the model capacity, training epochs and learning rate.
Passive Defense. Passive defense methods aim to reduce the impact of backdoor attacks during the inference process. For example, Qi et al. (2021a) proposed ONION, a method that detects and removes discrete words in sentences using perplexity and output probability outputted by the language model. Gao et al. (2021) proposed STRIP, which detects and removes backdoor samples by analyzing the model’s output. Jain et al. (2023) proposed a method that uses a generative model to interpret an adversarial instruction. Ideally, the generative model will accurately retain the natural instruction and may remove the malicious trigger in the instruction. Although the interpretation instruction works well in most cases, it may also cause model degradation.
@@ 7.3. Backdoor Defense
In our study, we focus on developing active defense methods against backdoor attacks. Our defense method leverages the ”early learning” phenomenon observed during the training of CLMs. Our proposed method not only showcases enhanced effectiveness but also exhibits a wider applicability scope when compared with previous defense methodologies.
@@ 8. Conclusion
While DeCE has shown promising results in defending against backdoor attacks, we want to optimize its hyper-parameters in the future, which can improve the defense quality against various attack strategies. Futhermore, we will explore more covert and complex poisoning attack methods to thoroughly evaluate the proposed defense mechanism’s performance in the real world. On the other hand, we would like to study its applicability in more areas of code intelligence and NLP, such as text classification, code summarization and other tasks.
@@ Acknowledgements.
This research/project is supported by the National Natural Science Foundation of China (No. 62372232), the Short-term Visiting Program of Nanjing University of Aeronautics and Astronautics for Ph.D. Students Abroad (No. 240602DF16), High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics, and the Collaborative Innovation Center of Novel Software Technology and Industrialization. T. Chen is partially supported by an oversea grant from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2022A03, KFKT2023A04). This research/project is supported by the National Research Foundation, under its Investigatorship Grant (NRF-NRFI08-2022-0002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.